OpenStudy (anonymous):

If the width of a rectangle is x + 2 and the length is twice as long, what is the area? A. 6x + 12 B. 2x² + 6x + 4 C. 2x² + 8x + 6 D. 2x² + 8x + 8

OpenStudy (twizttiez):

What do you think?

OpenStudy (princeharryyy):

area = length*breadth = (2*(x+2))*(x+2)

OpenStudy (princeharryyy):

2x² + 8x + 8

OpenStudy (anonymous):

okay what is the formula for area ? A=l*b given b= X +2 and l=2*b=2*(X+2) just substitute l and b in the eqn :) A would be in terms of x